OpenStudy (kayders1997):

The velocity function (in meters per second) is given for a particle moving along a line. Find a) the displacement and b) the distance traveled by the particle during the given time interval? V(t)=t^2-2t-8 1

OpenStudy (kayders1997):

@zepdrix

zepdrix (zepdrix):

The velocity function is the `derivative` of the displacement function. So then, we can find the displacement function by `anti-differentiation` of the velocity function, ya?

OpenStudy (kayders1997):

Yes

zepdrix (zepdrix):

\[\large\rm s(t)=\int\limits_1^6 v(t)dt\]

zepdrix (zepdrix):

\[\large\rm s(t)=\int\limits_1^6 t^2-2t-8~dt\]

zepdrix (zepdrix):

Apply `power rule for integration` to each term, ya? :)

OpenStudy (kayders1997):

Yes :O

OpenStudy (kayders1997):

Do you want me to do it? :O

zepdrix (zepdrix):

ya sure

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

1/3t^3-t^2-5t?

OpenStudy (kayders1997):

:o

zepdrix (zepdrix):

1/3t^3-t^2-8t, ya

OpenStudy (kayders1997):

Oops :D

zepdrix (zepdrix):

Then evaluate at the limits,\[\large\rm \left(\frac{1}{3}t^3-t^2-8t\right)_{t=6}-\left(\frac{1}{3}t^3-t^2-8t\right)_{t=1}\]

zepdrix (zepdrix):

\[\large\rm \left(\frac{1}{3}6^3-6^2-8(6)\right)-\left(\frac{1}{3}1^3-1^2-8(1)\right)\]

OpenStudy (kayders1997):

Woooooo :O

OpenStudy (kayders1997):

Okay just give me a second to figure this out

OpenStudy (kayders1997):

-10/3

OpenStudy (kayders1997):

@zepdrix

zepdrix (zepdrix):

yay good job

OpenStudy (kayders1997):

Now I have to do the other part :(

OpenStudy (kayders1997):

@zepdrix we have to do the next part!!!!

zepdrix (zepdrix):

soz

zepdrix (zepdrix):

\[\large\rm v(t)=t^2-2t-8\]We need to figure out when the velocity is zero, that tells us at which points the particle is changing directions.

zepdrix (zepdrix):

\[\large\rm 0=t^2-2t-8\]factor

OpenStudy (kayders1997):

(t-4)(t+2)

zepdrix (zepdrix):

Ok at time t=4 our velocity is zero, that is where it changes direction.

OpenStudy (kayders1997):

Oh you have to show that?

zepdrix (zepdrix):

So we have to do something fancy. We have to break up our integral.\[\large\rm \int\limits_1^6\quad=\quad \int\limits_1^4\quad+\quad\int\limits_4^6\]

OpenStudy (kayders1997):

Ugh I don't get doing that

zepdrix (zepdrix):

\[\large\rm d(t)=\left|\int\limits_1^4 t^2-2t-8~dt\right|+\left|\int\limits_4^6 t^2-2t-8~dt\right|\]

zepdrix (zepdrix):

no?

zepdrix (zepdrix):

here is an example